OpenStudy (xguardians):

Solve 11^(X+4) = 57 Please explain steps, ty!

OpenStudy (xguardians):

@Hero @SolomonZelman

OpenStudy (yanasidlinskiy):

Rearrange the equation by subtracting what is to the right of the greater than sign from both sides of the inequality : 11^(x+4)-(57)>0

OpenStudy (xguardians):

There is no greater than sign?

OpenStudy (solomonzelman):

\(\large\color{blue }{ \rm 11^{x+4}=57 }\) \(\large\color{blue }{ \rm (x+4)~log~11=log~57 }\) \(\large\color{blue }{ \rm (x+4)~=log~57~/~log~11 }\) \(\large\color{blue }{ \rm (x+4)~=log_{~~11}~57 }\) \(\large\color{blue }{ \rm x~=log_{~~11}~57~-4 }\)

OpenStudy (solomonzelman):

You just need to approximate the answer.

OpenStudy (xguardians):

@SolomonZelman Do you have a calculator on you? I don't have mine at the moment, left it at school, it would be great if you tell me answer :)?